• A 50.0 kg circus acrobat drops from a height of 2.0 m straight down onto a springboard with a force constant of 8.00 × 103 N/m. By what maximum distance does she compress the spring?

• Using energy considerations and assuming negligible air resistance, show that a rock thrown from a bridge 20.0 m above water with an initial speed of 15.0 m/s strikes the water with a speed of 24.8 m/s independent of the direction thrown.


1
Expert's answer
2020-07-06T17:11:58-0400

1.As "U_g=mgh,PE=\frac{1}{2}kx^2" then from "U_g=PE" hence "x=\sqrt\frac{2mgh}{k}=\sqrt\frac{2\sdot50\sdot9.8\sdot2}{8\sdot10^3}\approx0.5m"

2.As "mu^2\/2=mgh+mv^2\/2" hence "u=\sqrt{2gh+v^2}=\sqrt{2\sdot9.8\sdot20+15^2}\approx24.8m/c"